How do I find the equation of the line passing through the point (1,5) and parallel to 4x+7y=18?
Thank you for your time and consideration! I have a test on these tomorrow and I'm really confused.

Answer by scott8148(6628)

if the lines are parallel ; the equations are the same, except for the constant term (the 18)

to find the constant term for the new line ; plug the coordinates of the given point into the equation

4(1) + 7(5) = ? ___ 4 + 35 = ? ___ 39 = ?

the equation of the new line is ___ 4x + 7y = 39